# Intraoperative 2D/3D Registration via Spherical Similarity Learning and Inference-Time Differentiable Levenberg-Marquardt Optimization

### v1

@@ 1 Introduction
Intraoperative 2D/3D registration is a process aimed at aligning intraoperative 2D images, such as X-ray images, with corresponding preoperative CT scans. It is a crucial step in providing surgical planning guidance and navigation positioning for minimally invasive procedures. This process allows for the precise placement of implants or surgical instruments during spinal surgeries, such as percutaneous vertebroplasty and pedicle screw internal fixation [61].
@@
In conventional optimization-based registration methods [21, 47, 31, 14, 9, 37], diverse ray-tracing techniques are employed to produce simulated two-dimensional X-ray images, commonly referred to as digitally reconstructed radiographs (DRRs), from the 3D CT volume. These techniques simulate the attenuation of X-rays within the human body, providing a crucial component for accurate registration. The similarity between DRRs and X-rays is then evaluated by using some statistical-based similarity measures, ie., normalized cross-correlation (NCC) and mutual information (MI) or local feature representations [59]. Derivative-free optimization techniques, such as BOBYQA [49] and CMA-ES [27], are employed to explore the similarity function and identify both the minimum value and the corresponding patient pose within the solution space. Yet, such similarity functions are not globally convex in the search space, consequently presenting numerous local minima. Therefore, it is frequently employed alongside complex multi-resolution optimization strategies or necessitates a robust initialization.
@@
In literature, with the development of machine learning technology, several regression-based [34, 41, 32, 66] and landmark-based methods [12, 20, 2, 38] have been proposed, demonstrating promising performance. However, they suffer from either the vulnerability to minor image changes, leading to prediction instability, or demand an extensive amount of clinically unsatisfactory and nearly impractical well-annotated data for training. These limitations significantly hamper the clinical applicability of these methods. Recently, inspired by the traditional optimization-based methods, Gao et al. [15, 16] proposed a fully differentiable framework that approximates the geodesic between two points in a special Euclidean group in three dimensions (SE(3)) by using neural networks. This framework increases the capture range of registration while being more resilient to changes in local features. This constrained similarity learning approach greatly mitigates disturbances caused by large offsets and image noise, demonstrating robustness. Besides, an initial study [7] attempted to use a correlation-driven method to decompose local and global features during the similarity learning process. However, existing similarity learning-based frameworks still have several shortcomings. We observed that current methods opt to approximate manifold in Riemannian geometry within Euclidean space. However, Riemannian manifolds exhibit intricate local geometric structures. Approximating using Euclidean space leads to an inaccurate portrayal of the manifold’s local structure [67]. Because distances in feature space ought to be calculated along manifold directions, and the Euclidean distance serves as merely a suitable approximation for small distances. And we could not determine the precise distance metric in the latent space that is best suited for approximating geodesics on the special Euclidean group. Moreover, the convergence process of existing methods tends to be relatively protracted (see results in Tab. 2).
@@
In this work, we propose a more reasonable paradigm for the similarity learning-based 2D/3D registration. First, to mitigate the inaccuracies in local geometric descriptions when approximating Riemannian geodesic distance in Euclidean space, we explore alternative metric approximations in non-Euclidean embedding spaces and extend them to special orthogonal groups with bi-invariance. This means that the distance between two elements remains unchanged under both left and right multiplications, whereas SE(3) is only left-invariant. As a result, the landscape of the learned similarity metric becomes unambiguous, more contrastive, and smoother, enabling a broader capture range while preventing premature convergence. During experiments, we found that the similarity measure acquired through similarity learning in hyperspherical space yields the most effective registration outcomes, rectifying the inherent ambiguity of single-view 2D/3D registration. We introduce a differentiable Levenberg-Marquardt (LM) optimization strategy as an alternative to the gradient descent method to accelerate the convergence of registration. By conducting studies on pelvic data, spine data, and clinical data on endovascular thrombectomy for acute ischemic stroke, we demonstrate the superiority of our approach in the context of intraoperative CT to X-ray registration. Our method achieves state-of-the-art performance across multiple datasets, including pelvic and spine datasets, as well as clinical data for cerebral endovascular therapy, demonstrating its robustness and effectiveness in diverse 2D/3D registration scenarios. It should be noted that, compared with single-view registration, multi-view or biplanar 2D fluoroscopic registration to a 3D CT volume provides better global convexity and benefits from well-established solvers [40, 69, 35]. However, it requires acquiring additional fluoroscopic images, leading to higher radiation exposure that may harm patient health. In this work, we focus on single-view 2D/3D registration, aiming to balance accuracy with clinical safety.
@@ 2.1 Learning-based 2D/3D Registration
Learning-based methods in 2D/3D registration are primarily categorized into regression-based and landmark-based approaches. Regression-based methods [26, 32, 42, 66], directly estimate human pose from 3D CT volumes or paired 2D X-ray/DRR images. The primary advantage of this approach lies in its ability to achieve real-time posture estimation, making it particularly suitable for time-sensitive scenarios, e.g. intraoperative navigation and robotic-assisted surgery. However, its major limitation is relatively lower accuracy compared to landmark-based methods, which may impact precision in high-stakes medical scenarios requiring meticulous anatomical alignment. The landmark-based approach [12, 20, 35, 54, 55] formulates this task as a well-established perspective-n-point (PnP) problem [33], which determines the patient’s actual posture by estimating the spatial transformation of corresponding landmarks between the reference and moving images. Additionally, a point-to-plane correspondence solver [30, 53, 62, 29] has been proposed as an alternative approach to address this problem. While these methods improve alignment accuracy, it is highly dependent on the precision and visibility of landmarks [59], making them susceptible to noise, occlusions, and variations in image resolution. Furthermore, to enhance robustness, sophisticated techniques such as RANSAC [13] are often required, leading to increased computational cost and processing time. Recently, some studies [8, 64, 18, 11, 19] have adopted a two-stage approach to achieve robust and high-precision registration. This method first utilizes a regressor to initialize the pose, followed by fine-tuning through either differentiable similarity learning-based [17] or non-derivative optimization [34]. Literature [18] indicates that this approach can achieve sub-millimeter registration accuracy within a few seconds in patient-specific scenarios.
@@ 2.2 Similarity Learning for Image Registration
Image registration is often regarded as a task without a definitive gold standard, making the evaluation of image similarity a critical challenge [37]. Traditional similarity metrics, such as sum of squared differences (SSD), NCC, and MI, exhibits convexity around the optimum but non-convexity elsewhere, which can trap the estimates into undesirable local minima. Similarity learning enables the generation of a globally convex and smooth optimization landscape by allowing neural networks to learn feature representations of images [7, 50, 23]. The resulting similarity metric can serve as a differentiable approximation of traditional metrics [51], a deep distance metric capturing the inherent geometric structure of the image [24, 57], or a deep neighborhood self-similarity measure that enhances robustness in registration tasks [43]. Gu et al.[25] first introduced similarity learning into the field of 2D/3D registration by proposing a method to learn a convex similarity metric through the estimation of the Riemannian pose gradient. Building on this foundation, Gao et al. [15, 16] extended the approach to a fully differentiable framework, while Chen et al. [8] designed a parameter-specific pose initialization strategy to enhance its performance. In contrast to these existing methods, our approach considers the approximation of the distance gradient on the Riemannian manifold in a non-Euclidean space and further extends the learning strategy from SE(3) to SO(4). This extension ensures that the learned deep metric inherits the bi-invariant properties of SO(4), leading to improved stability and consistency in geometric optimization for 2D/3D registration.
@@ 3.2 Network Architecture
The overall framework of our proposed method is shown in Fig. 1. Let $I$ be the set of X-ray images and $V$ be the set of CT volumes. Similar to many recent approaches [8, 18, 64], we employ a regressor $\mathcal{R}(\cdot)$ to initialize the pose. Our framework adapts the choice of the regressor based on the application scenario. In patient-agnostic scenarios, the regressor takes both $V$ and $I$ as inputs. However, in patient-specific scenarios, where training is performed exclusively on a subject’s CT and corresponding X-ray images, the regressor operates solely on $I$. This is because patient-specific registration does not require learning the anatomical variability present in the CT volume across different individuals, allowing the model to focus on refining pose estimation based on the subject’s X-ray features. The regression output $\theta_{ini}$ is parameterized in $\mathfrak{se}(3)$, which is isomorphic to $\mathbb{R}^{3}\times\mathbb{R}^{3}$ and jointly represents an axis-angle rotation and a translation vector [46]. This parameterization has been shown to achieve the highest accuracy in 2D/3D registration [18]. Next, we map $\theta_{ini}$ to the SE(3) to obtain $\textbf{T}_{ini}$, and then use the projector $\mathcal{P}(\cdot)$ to generate the DRR image $I_{ini}$ at the initial pose defined by $\textbf{T}_{ini}$.
Previous study [16] has shown that passing $V$ through a residual 3D CNN before projection can enhance registration accuracy by introducing a learnable parameter module in the 3D domain. We hypothesize that this CNN module captures intensity information to compensate for the domain gap between the projected DRRs and real X-ray images. However, while CNNs can achieve accurate rigid registration, their design does not inherently exploit the symmetry of rigid motion. Convolutional kernels are naturally equivariant to translation, meaning their output shifts with the input, but they lack equivariance to rotation. As a result, CNNs may extract different features for the same object in different poses, making it challenging to learn features that remain consistent under rigid transformations. To address the limitation above, we employ a Steerable Equivariant CNNs (E-CNNs) [45, 1] $\Psi(\cdot)$ to extract robust SE(3)-equivariant feature maps. A neural network is considered SE(3)-equivariant if, under a rigid body transformation T, it satisfies the following condition for an input volume $V$:
@@ 3.3 Spherical Similarity Learning
where $\mathrm{EXP}(\cdot)$ denotes the spherical exponential map [63]. This map is defined such that any vector $\phi\in\mathbb{R}^{d}$ is mapped to a point $\Phi$ on the unit sphere $S^{d}\subset\mathbb{R}^{d+1}$. We then construct a vector $\bar{\phi}=[\phi,1]\in T_{\textbf{N}}\mathbb{S}^{d}$ in the tangent space of the north pole $\textbf{N}=[0,0,...,0,1]\in\mathbb{S}^{d}$ and it can be expressed as:
@@
In this work, we adopt the double-backward strategy [16] to approximate the gradient of the network’s loss with the gradient of a geodesic distance on the pose manifold. This approach has proven effective in mitigating sensitivity to absolute loss scaling, as it prioritizes preserving the local geometric structure of the manifold rather than relying solely on Euclidean differences. Specifically, let $\theta\in\mathfrak{se}(3)$ denote the pose of the camera. We denote by $\theta_{gt}$ the corresponding ground-truth pose. Our network produces a similarity score $\mathcal{L}_{\mathrm{net}}=\varepsilon$, whose gradient with respect to $\theta$ is $\nabla_{\theta}\,\mathcal{L}_{\mathrm{net}}$.
Geodesic distance for gradient approximation. We introduce a geodesic loss $\mathcal{L}_{\mathrm{geo}}(\theta_{A},\theta_{B})$ to measure the distance between two poses, e.g. $\theta_{A}$ and $\theta_{B}$, on Riemannian manifold. One common choice is to measure the pose error on $\mathfrak{se}(3)$ [28], and the geodesic distance on $\mathfrak{se}(3)$ is given by
@@
where $\mathcal{LOG}:\text{SE}(3)\rightarrow\mathfrak{se}(3)$ and $\mathcal{EXP}:\mathfrak{se}(3)\rightarrow\text{SE}(3)$ are the matrix logarithm and exponential mapping between $\mathfrak{se}(3)$ and $\text{SE}(3)$. This geodesic definition captures the angular distance on the manifold, rather than an absolute Euclidean difference [36, 52]. However, since the geodesic distance on SE(3) is only left-invariant, the learned similarity function may exhibit ambiguity and inconsistency in its optimization landscape. To address this issue, we explore approximating the gradient within the bi-invariant structure of SO(4), which provides a more stable and geometrically consistent formulation. Following established approaches [39, 58], the mapping from $\theta\in\mathfrak{se}(3)\rightarrow\textbf{T}\in\mathrm{SE}(3)$ to $\textbf{H}\in\mathrm{SO}(4)$ is defined as:
@@
To avoid the impact of absolute loss scale, we follow the idea of previous works [16, 7] and approximate the gradient of $\mathcal{L}_{\mathrm{net}}$ by aligning it with the gradient of the geodesic distance. Concretely, we want: $\nabla_{\theta}\,\mathcal{L}_{\mathrm{net}}(\theta)\;\approx\;\nabla_{\theta}\,\mathcal{L}_{\mathrm{geo}}\!\bigl{(}\theta,\theta_{gt}\bigr{)}$. We thus define an approximate loss $\mathcal{L}$ that measures the mismatch between these two gradients:

### v3

# Intraoperative 2D/3D Registration via Spherical Similarity Learning and Differentiable Levenberg-Marquardt Optimization

@@ 1 Introduction
Intraoperative 2D/3D registration is a process aimed at aligning intraoperative 2D images, such as X-ray images, with corresponding preoperative CT scans. It is a crucial step in providing surgical planning guidance and navigation positioning for minimally invasive procedures. This process allows for the precise placement of implants or surgical instruments during spinal surgeries, such as percutaneous vertebroplasty and pedicle screw internal fixation [unberath2021impact].
@@
In conventional optimization-based registration methods [grupp2018patch, penney1998comparison, knaan2003effective, frysch2021novel, de20163d, markelj2012review], diverse ray-tracing techniques are employed to produce simulated two-dimensional X-ray images, commonly referred to as digitally reconstructed radiographs (DRRs), from the 3D CT volume. These techniques simulate the attenuation of X-rays within the human body, providing a crucial component for accurate registration. The similarity between DRRs and X-rays is then evaluated by using some statistical-based similarity measures, ie., normalized cross-correlation (NCC) and mutual information (MI) or local feature representations [toews2017phantomless]. Derivative-free optimization techniques, such as BOBYQA [powell2009bobyqa] and CMA-ES [hansen2001completely], are employed to explore the similarity function and identify both the minimum value and the corresponding patient pose within the solution space. Yet, such similarity functions are not globally convex in the search space, consequently presenting numerous local minima. Therefore, it is frequently employed alongside complex multi-resolution optimization strategies or necessitates a robust initialization.
@@
In literature, with the development of machine learning technology, several regression-based [li2025automatic, miao2016cnn, leroy2023structuregnet, zhao2024automatic] and landmark-based methods [esteban2019towards, grimm2021pose, brandstatter2024rigid, markova2022global] have been proposed, demonstrating promising performance. However, they suffer from either the vulnerability to minor image changes, leading to prediction instability, or demand an extensive amount of clinically unsatisfactory and nearly impractical well-annotated data for training. These limitations significantly hamper the clinical applicability of these methods. Recently, inspired by the traditional optimization-based methods, Gao et al. [gao2020generalizing, gao2023fully] proposed a fully differentiable framework that approximates the geodesic between two points in a special Euclidean group in three dimensions (SE(3)) by using neural networks. This framework increases the capture range of registration while being more resilient to changes in local features. This constrained similarity learning approach greatly mitigates disturbances caused by large offsets and image noise, demonstrating robustness. Besides, an initial study [chen2024fully] attempted to use a correlation-driven method to decompose local and global features during the similarity learning process. However, existing similarity learning-based frameworks still have several shortcomings. We observed that current methods opt to approximate manifold in Riemannian geometry within Euclidean space. However, Riemannian manifolds exhibit intricate local geometric structures. Approximating using Euclidean space leads to an inaccurate portrayal of the manifold’s local structure [zhao2023spherical]. Because distances in feature space ought to be calculated along manifold directions, and the Euclidean distance serves as merely a suitable approximation for small distances. And we could not determine the precise distance metric in the latent space that is best suited for approximating geodesics on the special Euclidean group. Moreover, the convergence process of existing methods tends to be relatively protracted (see results in Tab. 2).
@@
In this work, we propose a more reasonable paradigm for the similarity learning-based 2D/3D registration. First, to mitigate the inaccuracies in local geometric descriptions when approximating Riemannian geodesic distance in Euclidean space, we explore alternative metric approximations in non-Euclidean embedding spaces and extend them to special orthogonal groups with bi-invariance. This means that the distance between two elements remains unchanged under both left and right multiplications, whereas SE(3) is only left-invariant. As a result, the landscape of the learned similarity metric becomes unambiguous, more contrastive, and smoother, enabling a broader capture range while preventing premature convergence. During experiments, we found that the similarity measure acquired through similarity learning in hyperspherical space yields the most effective registration outcomes, rectifying the inherent ambiguity of single-view 2D/3D registration. We introduce a differentiable Levenberg-Marquardt (LM) optimization strategy as an alternative to the gradient descent method to accelerate the convergence of registration. By conducting studies on pelvic data, spine data, and clinical data on endovascular thrombectomy for acute ischemic stroke, we demonstrate the superiority of our approach in the context of intraoperative CT to X-ray registration. Our method achieves state-of-the-art performance across multiple datasets, including pelvic and spine datasets, as well as clinical data for cerebral endovascular therapy, demonstrating its robustness and effectiveness in diverse 2D/3D registration scenarios. It should be noted that, compared with single-view registration, multi-view or biplanar 2D fluoroscopic registration to a 3D CT volume provides better global convexity and benefits from well-established solvers [mi2023sgreg, zollei20012d, liao2019multiview]. However, it requires acquiring additional fluoroscopic images, leading to higher radiation exposure that may harm patient health. In this work, we focus on single-view 2D/3D registration, aiming to balance accuracy with clinical safety.
@@ 2.1 Learning-based 2D/3D Registration
Learning-based methods in 2D/3D registration are primarily categorized into regression-based and landmark-based approaches. Regression-based methods [guo2021end, leroy2023structuregnet, miao2016real, zhao2024automatic], directly estimate human pose from 3D CT volumes or paired 2D X-ray/DRR images. The primary advantage of this approach lies in its ability to achieve real-time posture estimation, making it particularly suitable for time-sensitive scenarios, e.g. intraoperative navigation and robotic-assisted surgery. However, its major limitation is relatively lower accuracy compared to landmark-based methods, which may impact precision in high-stakes medical scenarios requiring meticulous anatomical alignment. The landmark-based approach [esteban2019towards, grimm2021pose, liao2019multiview, shrestha2023x, shrestha2024rayemb] formulates this task as a well-established perspective-n-point (PnP) problem [li2012robust], which determines the patient’s actual posture by estimating the spatial transformation of corresponding landmarks between the reference and moving images. Additionally, a point-to-plane correspondence solver [jaganathan2023self, schaffert2020learning, wang2017dynamic, jaganathan2021deep] has been proposed as an alternative approach to address this problem. While these methods improve alignment accuracy, it is highly dependent on the precision and visibility of landmarks [toews2017phantomless], making them susceptible to noise, occlusions, and variations in image resolution. Furthermore, to enhance robustness, sophisticated techniques such as RANSAC [fischler1981random] are often required, leading to increased computational cost and processing time. Recently, some studies [chen2024embedded, zhang2023patient, gopalakrishnan2024intraoperative, downs2025improving, gopalakrishnan2025rapid] have adopted a two-stage approach to achieve robust and high-precision registration. This category of methods first utilizes a regressor to initialize the pose, followed by fine-tuning through either differentiable [gopalakrishnan2022fast] or non-derivative optimization [li2025automatic]. Literature [gopalakrishnan2024intraoperative] indicates that this approach can achieve sub-millimeter registration accuracy within a few seconds in patient-specific scenarios.
@@ 2.2 Similarity Learning for Image Registration
Image registration is often regarded as a task without a definitive gold standard, making the evaluation of image similarity a critical challenge [markelj2012review]. Traditional similarity metrics, such as sum of squared differences (SSD), NCC, and MI, exhibits convexity around the optimum but non-convexity elsewhere, which can trap the estimates into undesirable local minima. Similarity learning enables the generation of a globally convex and smooth optimization landscape by allowing neural networks to learn feature representations of images [chen2024fully, qin2019unsupervised, grzech2022variational]. The resulting similarity metric can serve as a differentiable approximation of traditional metrics [ronchetti2023disa], a deep distance metric capturing the inherent geometric structure of the image [grzech2024unsupervised, sideri2023mad], or a deep neighborhood self-similarity measure that enhances robustness in registration tasks [mok2024modality]. Gu et al.[gu2020extended] first introduced similarity learning into the field of 2D/3D registration by proposing a method to learn a convex similarity metric through the estimation of the Riemannian pose gradient. Building on this foundation, Gao et al. [gao2020generalizing, gao2023fully] extended the approach to a fully differentiable framework, while Chen et al. [chen2024embedded] designed a parameter-specific pose initialization strategy to enhance its performance. In contrast to these existing methods, our approach considers the approximation of the distance gradient on the Riemannian manifold in a non-Euclidean space and further extends the learning strategy from SE(3) to SO(4). This extension ensures that the learned deep metric inherits the bi-invariant properties of SO(4), leading to improved stability and consistency in geometric optimization for 2D/3D registration.
@@ 3.2 Network Architecture
The overall framework of our proposed method is shown in Fig. 1. Let $I$ be the set of X-ray images and $V$ be the set of CT volumes. Similar to many recent approaches [chen2024embedded, gopalakrishnan2024intraoperative, zhang2023patient], we employ a regressor $\mathcal{R}(\cdot)$ to initialize the pose. Our framework adapts the choice of the regressor based on the application scenario. In patient-agnostic scenarios, the regressor takes both $V$ and $I$ as inputs. However, in patient-specific scenarios, where training is performed exclusively on a subject’s CT and corresponding X-ray images, the regressor operates solely on $I$. This is because patient-specific registration does not require learning the anatomical variability present in the CT volume across different individuals, allowing the model to focus on refining pose estimation based on the subject’s X-ray features. The regression output $\theta_{ini}$ is parameterized in $\mathfrak{se}(3)$, which is isomorphic to $\mathbb{R}^{3}\times\mathbb{R}^{3}$ and jointly represents an axis-angle rotation and a translation vector [murray2017mathematical]. This parameterization has been shown to achieve the highest accuracy in 2D/3D registration [gopalakrishnan2024intraoperative]. Next, we map $\theta_{ini}$ to the SE(3) to obtain $\textbf{T}_{ini}$, and then use the projector $\mathcal{P}(\cdot)$ to generate the DRR image $I_{ini}$ at the initial pose defined by $\textbf{T}_{ini}$.
Previous study [gao2023fully] has shown that passing $V$ through a residual 3D CNN before projection can enhance registration accuracy by introducing a learnable parameter module in the 3D domain. We hypothesize that this CNN module captures intensity information to compensate for the domain gap between the projected DRRs and real X-ray images. However, while CNNs can achieve accurate rigid registration, their design does not inherently exploit the symmetry of rigid motion. Convolutional kernels are naturally equivariant to translation, meaning their output shifts with the input, but they lack equivariance to rotation. As a result, CNNs may extract different features for the same object in different poses, making it challenging to learn features that remain consistent under rigid transformations. To address the limitation above, we employ a Steerable Equivariant CNNs (E-CNNs) [moyer2021equivariant, billot2024se] $\Psi(\cdot)$ to extract robust SE(3)-equivariant feature maps. A neural network is considered SE(3)-equivariant if, under a rigid body transformation T, it satisfies the following condition for an input volume $V$:
@@ 3.3 Spherical Similarity Learning
where $\mathrm{EXP}(\cdot)$ denotes the spherical exponential map [wilson2014spherical]. This map is defined such that any vector $\phi\in\mathbb{R}^{d}$ is mapped to a point $\Phi$ on the unit sphere $S^{d}\subset\mathbb{R}^{d+1}$. We then construct a vector $\bar{\phi}=[\phi,1]\in T_{\textbf{N}}\mathbb{S}^{d}$ in the tangent space of the north pole $\textbf{N}=[0,0,...,0,1]\in\mathbb{S}^{d}$ and it can be expressed as:
@@
In this work, we adopt the double-backward strategy [gao2023fully] to approximate the gradient of the network’s loss with the gradient of a geodesic distance on the pose manifold. This approach has proven effective in mitigating sensitivity to absolute loss scaling, as it prioritizes preserving the local geometric structure of the manifold rather than relying solely on Euclidean differences. Specifically, let $\theta\in\mathfrak{se}(3)$ denote the pose of the camera. We denote by $\theta_{gt}$ the corresponding ground-truth pose. Our network produces a similarity score $\mathcal{L}_{\mathrm{net}}=\varepsilon$, whose gradient with respect to $\theta$ is $\nabla_{\theta}\,\mathcal{L}_{\mathrm{net}}$.
Geodesic distance for gradient approximation. We introduce a geodesic loss $\mathcal{L}_{\mathrm{geo}}(\theta_{A},\theta_{B})$ to measure the distance between two poses, e.g. $\theta_{A}$ and $\theta_{B}$, on Riemannian manifold. One common choice is to measure the pose error on $\mathfrak{se}(3)$ [hou2018computing], and the geodesic distance on $\mathfrak{se}(3)$ is given by
@@
where $\mathcal{LOG}:\text{SE}(3)\rightarrow\mathfrak{se}(3)$ and $\mathcal{EXP}:\mathfrak{se}(3)\rightarrow\text{SE}(3)$ are the matrix logarithm and exponential mapping between $\mathfrak{se}(3)$ and $\text{SE}(3)$. This geodesic definition captures the angular distance on the manifold, rather than an absolute Euclidean difference [mahendran20173d, salehi2018real]. However, since the geodesic distance on SE(3) is only left-invariant, the learned similarity function may exhibit ambiguity and inconsistency in its optimization landscape. To address this issue, we explore approximating the gradient within the bi-invariant structure of SO(4), which provides a more stable and geometrically consistent formulation. Following established approaches [meng2025parametric, thomas2014approaching], the mapping from $\theta\in\mathfrak{se}(3)\rightarrow\textbf{T}\in\mathrm{SE}(3)$ to $\textbf{H}\in\mathrm{SO}(4)$ is defined as:
@@
To avoid the impact of absolute loss scale, we follow the idea of previous works [gao2023fully, chen2024fully] and approximate the gradient of $\mathcal{L}_{\mathrm{net}}$ by aligning it with the gradient of the geodesic distance. Concretely, we want: $\nabla_{\theta}\,\mathcal{L}_{\mathrm{net}}(\theta)\;\approx\;\nabla_{\theta}\,\mathcal{L}_{\mathrm{geo}}\!\bigl(\theta,\theta_{gt}\bigr)$. We thus define an approximate loss $\mathcal{L}$ that measures the mismatch between these two gradients: